# ODEStream: A Buffer-Free Online Learning Framework with ODE-based Adaptor for Streaming Time Series Forecasting

### v1

@@ Abstract
Addressing the challenges of irregularity and concept drift in streaming time series is crucial in real-world predictive modelling. Previous studies in time series continual learning often propose models that require buffering of long sequences, potentially restricting the responsiveness of the inference system. Moreover, these models are typically designed for regularly sampled data, an unrealistic assumption in real-world scenarios. This paper introduces ODEStream, a novel buffer-free continual learning framework that incorporates a temporal isolation layer that integrates temporal dependencies within the data. Simultaneously, it leverages the capability of neural ordinary differential equations to process irregular sequences and generate a continuous data representation, enabling seamless adaptation to changing dynamics in a data streaming scenario. Our approach focuses on learning how the dynamics and distribution of historical data change with time, facilitating direct processing of streaming sequences. Evaluations on benchmark real-world datasets demonstrate that ODEStream outperforms the state-of-the-art online learning and streaming analysis baselines, providing accurate predictions over extended periods while minimising performance degradation over time by learning how the sequence dynamics change. Our code is available at: https://anonymous.4open.science/r/ODEStream-2BAC.
@@ 1 Introduction
Handling streaming analysis requires an effective continual online learning approach. Recently few studies have addressed streaming analysis for regression and forecasting (Cossu et al., 2021), by relying on buffer memory to apply a replay method for continual learning (Matteoni et al., 2022; Kwon et al., 2021; Kiyasseh et al., 2021; Xiao et al., 2022; Chen et al., 2021). Using the replay method, a training set initialises the model’s parameters, and a selected subset of this data, as well as future incoming data samples, is saved to be used later when the model is retrained (Hoi et al., 2021; Fekri et al., 2021; He, 2021). This process focuses on avoiding catastrophic forgetting issues by continuously updates the model knowledge while retraining on both incoming and historical data. However, the effectiveness of this method is questionable when applied to time series data, as it primarily focuses on preventing the loss of previously learned information. This may lead to an insufficient learning on new dependencies evolving in the data over time (Cossu et al., 2021; Pham et al., 2023). Additionally, these reply-based frameworks involve many complexities as it needs to carefully determine what data samples are important to store in the buffer, when to update the importance weights, and how to allocate memory and resources effectively. These complexities underscore the need for a more streamlined approach to managing streaming time series data. Furthermore, catastrophic forgetting impact differs when dealing with time series data and predictive modelling compared to non-temporal data and classification tasks where class and task-incremental learning are common scenarios (Ao & Fayek, 2023; Pham et al., 2023). For such tasks the model objective is to learn new tasks or classes where it is important to make sure that the old tasks and classes are not forgotten. In the context of developing a lifelong learning model for streaming time series data, a relevant consideration arises related to the extent to which we should prioritise preserving historical information that represents data that has evolved and changed (Cossu et al., 2021; Pham et al., 2023).
@@
Streaming time series forecasting faces significant challenges, including (1) Concept Drift Phenomenon, where underlying patterns or relationships within the data change over time, posing a substantial challenge for predictive modelling (Fekri et al., 2021). This issue extends beyond online learning models to impact batch learning approaches, as models initially proven accurate can quickly become obsolete as the data continues to evolve. (2) Temporal Irregularity which is common in streaming data, especially in real-world applications where the data lacks a specific time frame. (3) Dynamic Preservation where emphasises the importance of striking the right balance between old and new data. This balance is critical to preventing biases towards outdated data while effectively adapting to the dynamic nature of streaming time series data. Building a robust framework that can adeptly address these challenges is essential for maintaining model accuracy over time.
@@
This paper aims to overcome the restrictive boundaries of streaming time series data processing by presenting a novel model (ODEStream) for online learning. Our focus lies in the model’s ability to adapt to changes without a complicated framework while maintaining good performance over time. Specifically, we investigate how the novel neural ODE methods can be leveraged for memory-free online forecasting on a streaming time series. This paper includes the following contributions:
  • Unique Perspective: We redefine the context of continuous time series forecasting where our approach revolves around a unique perspective that emphasises the importance of learning the dynamic patterns and distribution of the time series data. By prioritising these aspects, our model shows a higher potential compared to traditional methods that focus on preserving historical data samples.
@@
  • Utilising Neural ODE: To the best of our knowledge, our work represents the first attempt to leverage novel neural Ordinary Differential Equation (ODE) models in the context of continuous learning and online forecasting. Due to their ability to dynamically adapt and provide a continuous hidden state, ODEs offer a more flexible approach compared to traditional discrete-time models, especially in handling irregularly sampled data.
@@
  • Streamlined Buffer-Free Framework: We introduce ODEStream, a buffer-free continual learning framework designed to have a straightforward and efficient training process. Unlike other frameworks that require intricate decision-making for upcoming data samples, our approach eliminates the need to define thresholds, trigger values, or buffering samples to regulate the algorithm’s flow and parameter updates. This simplification not only enhances the model’s ease of use but also eliminates the burden of complex decision requirements associated with existing frameworks.
@@
  • Comprehensive Evaluation: We evaluate ODEStream against the state-of-the-art continual learning models and show that ODEStream prevents performance from being affected by irregularity and effectively adapts to the drift that occurs in the data as well as maintaining a stable performance over very long streaming periods.
@@ 2.2 Time Series Streaming Data
In general, online time series forecasting has not been widely studied using deep learning models. Most of the traditional works have focused on machine learning models (Jiménez-Herrera et al., 2023; Li et al., 2019; Chen et al., 2020; Pham et al., 2023; Melgar-García et al., 2023; Cossu et al., 2021; Fekri et al., 2021; Ao & Fayek, 2023), or statistical forecasting techniques (Alberg & Last, 2018), either using online-offline learning (Chen et al., 2020; Pham et al., 2023; Melgar-García et al., 2023), or by applying incremental learning (Alberg & Last, 2018; Li et al., 2019). However, using machine learning is not sufficient for multivariate real-world data, where the temporal dependence and pattern are complex and hard to discover. Recent work has focused on online time series forecasting using deep learning. The work in (Wambura et al., 2020) introduced a model named OFAT, which stands for One Sketch Fits All-Time Series. The method was designed for long-range forecasting in feature-evolving data streams, which involve data that changes over time and has varying features. The model employs deep neural networks to learn from a small sample of data streams and subsequently applies the learned patterns to new data streams. For that work, CNN is used to extract features from the input data, while the LSTM is used to predict future values of the output data. In another work (Pham et al., 2023), the authors introduced a novel framework for online time series forecasting that combines the strengths of deep neural networks and Complementary Learning Systems (CLS). The framework, named FSNet, consists of two components: a slowly-learned backbone and a fast adapter. The backbone is responsible for learning long-term temporal dependencies, while the adapter is responsible for learning short-term changes. The two components are dynamically balanced to ensure that the model can adapt to both new and recurring patterns. When the data distribution changes, the adapter is updated to learn the new patterns. When the data distribution is stable, the backbone is updated to improve the accuracy of the prediction, while memory is used to store the historical data that the model has learned. Nevertheless, these works continue to face challenges when dealing with irregular samples. More Recent studies have focused on the temporal distribution shift problem. In (Du et al., 2021), the concept of Temporal Covariate Shift (TCS) was introduced, and the AdaRNN framework was proposed. AdaRNN combines Temporal Distribution Characterisation (TDC) and Temporal Distribution Matching (TDM) to adaptively manage distribution changes. While the experimental results show that AdaRNN significantly improves forecasting accuracy, it relies heavily on correctly segmenting the data into diverse periods. Bai et al. (2022) introduced the DRAIN framework, which also focuses on adapting to changing data distributions over time by addressing the challenge of temporal domain generalization. DRAIN uses a Bayesian approach to jointly model data and neural network dynamics. It employs a recurrent graph generation scenario to encode and decode dynamic neural networks, capturing temporal drifts in both data and model parameters. A major drawback of the model is its requirement to encode and decode the entire network parameters, which adds significant complexity to the model, making it computationally intensive.
@@ 2.3 Neural ODEs
Neural Ordinary Differential Equations (Neural ODEs) (Chen et al., 2018) represent a significant advancement in deep learning by integrating differential equations and neural networks. They provide a flexible framework for modelling complex dynamical systems, treating neural networks as continuous processes. Recent research in this field has made remarkable progress, showcasing the effectiveness of neural ODEs in various applications, especially for modelling irregularly sampled time series and partially observed data such as ODE-RNN, latent-ODE, NJ-ODE and GRU-ODE models (Rubanova et al., 2019; Herrera et al., 2020; De Brouwer et al., 2019). These models leverage the ODE’s capability to provide continuous hidden states, unlike the traditional discrete approach used in traditional sequence models such as RNN.
In recent years, several models based on differential equations (DE) have been introduced to address different issues, including modifying the hidden trajectories (Kidger et al., 2020; Morrill et al., 2021; Schirmer et al., 2022) or reducing computational overhead (Habiba & Pearlmutter, 2020). However, these models have typically been evaluated and applied in batch learning and fixed datasets; hence, we do not consider them in this work. Despite the progress in applying neural ODEs to various domains, a critical gap exists in their application to streaming time series data. Existing works primarily focus on batch learning and fixed datasets, lacking exploration of the dynamic and evolving nature of streaming data. The intricacies of adapting neural ODEs to real-time processing of continuously changing data remain unexplored. Our research aims to bridge this gap by introducing ODEStream, a novel framework for continual learning in streaming time series, effectively utilising the strengths of neural ODEs in this challenging context. Moreover, while a recent work (Giannone et al., 2020) proposed real-time classification from event-camera streams using neural ODEs (INODE), its focus on high-frequency event data and real-time classification does not comprehensively address the broader spectrum of challenges associated with streaming time series prediction, which we aim to tackle in this study.
@@ 3 Preliminaries
In this section, we outline the essential principles of ODE-based models and the concept of ODE solvers. the theory of the neural ODE and ODE Solver, and how they represent the irregular time series
@@ 4 Methodology
Problem statement: Let us consider a dynamic environment where streaming time series data is continuously generated as $X=\{x_{1},x_{2},\dots,x_{t},\dots\}$ with observations arriving in sequence over time. The goal is to develop a continual learning system for time series forecasting that adapts to the evolving nature of the data stream. Each observation $x_{t}$ represents the input at time $t$ for univariate data, $x_{t}$ represents a vector of inputs at time $t$ for multivariate data, such as $x_{t}=[x_{t1},x_{t2},...,x_{td}]$, where $x_{td}$ is the value of the $d^{th}$ feature at time $t$. The objective is to predict the future values of the time series. The challenge lies in the continual learning setting, where the model must adapt and learn from the continuously incoming data without access to past observations.
Figure 1: The ODEStream Framework leverages Neural ODEs to encode prior knowledge in the initial training phase. Subsequently, the learned model parameters are transferred and complemented by the temporal isolation layer for continuous online learning. In real-time, streaming data samples are used to predict future values, and concurrently, the model dynamically updates its parameters to adapt to newly observed samples.
@@ 4.1 Neural ODEs for Continual Learning
Our goal is to design a training method for memory-free continual learning. As mentioned before, in the context of online learning, we want our model to keep learning and adapting as new data arrives while avoiding the problem of forgetting what it has previously learned. Typically, complex model setups with buffers have been employed to store a subset of past data to address this problem. As previously discussed, studies in time series continual learning have shown that memory-based approaches yield the best performance, especially in classification tasks. The models in this framework continuously monitor incoming data and retrain on both old and new samples to maintain accuracy in a dynamic environment, as demonstrated by Fekri et al. (2021). However, when these models are applied to a continuous data stream, particularly one with gradually shifting data distributions, their effectiveness becomes uncertain. This uncertainty arises because the model only has one opportunity to train on each data sequence, unlike batch learning. Furthermore, in the presence of evolving temporal patterns, it remains unclear how the information stored in the buffer memory can still contribute to future prediction tasks.
@@ 4.2 Overview of ODEStream
We proposed a streamlined model for continual learning and forecasting of time series, consisting of two main phases. $(i)$ an offline learning phase, commonly referred to as model warm-up, designed to initialise the model parameters. $(ii)$ online learning and forecasting, which generate predictions based on new sequence flow while adopting and learning from the new changes in the data. The first phase is instrumental in obtaining prior knowledge of the available time sequences. For this part, the model utilises variational autoencoders to capture the distribution, dynamics, and temporal dependencies of historical data, while an ODE is employed to model evolving dynamics. The trained model is then transitioned to online learning, functioning as a dynamic adaptive network for continuous analysis of streaming sequences. In this primary phase, we leverage pre-trained NeuralODE weights. The model dynamically adapts to changes in the data distribution by continuously updating its understanding of the underlying dynamics. As in Figure 1, it’s noteworthy that the model framework acks external memory for storing samples lacks a decision node and threshold values to control the flow of the learning process. In the following section, we provide a detailed description of the model framework.
4.3 Obtaining Prior Knowledge.

### v2

@@ Abstract
Addressing the challenges of irregularity and concept drift in streaming time series is crucial for real-world predictive modelling. Previous studies in time series continual learning often propose models that require buffering long sequences, potentially restricting the responsiveness of the inference system. Moreover, these models are typically designed for regularly sampled data, an unrealistic assumption in real-world scenarios. This paper introduces ODEStream, a novel buffer-free continual learning framework that incorporates a temporal isolation layer to capture temporal dependencies within the data. Simultaneously, it leverages the capability of neural ordinary differential equations to process irregular sequences and generate a continuous data representation, enabling seamless adaptation to changing dynamics in a data streaming scenario. Our approach focuses on learning how the dynamics and distribution of historical data change over time, facilitating direct processing of streaming sequences. Evaluations on benchmark real-world datasets demonstrate that ODEStream outperforms the state-of-the-art online learning and streaming analysis baseline models, providing accurate predictions over extended periods while minimising performance degradation over time by learning how the sequence dynamics change. The implementation of ODEStream is available at: https://github.com/FtoonAbushaqra/ODEStream.git.
@@ 1 Introduction
Handling streaming analysis requires an effective continual online learning approach. Recently few studies have addressed streaming analysis for regression and forecasting (Cossu et al., 2021), by relying on buffer memory to apply a replay method for continual learning (Matteoni et al., 2022; Kwon et al., 2021; Kiyasseh et al., 2021; Xiao et al., 2022; Chen et al., 2021). Using the replay method, a training set initialises the model’s parameters, and a selected subset of this data, as well as future incoming data samples, is saved to be used later when the model is retrained (Hoi et al., 2021; Fekri et al., 2021; He, 2021). This process focuses on avoiding catastrophic forgetting issues by continuously updates the model knowledge while retraining on both incoming and historical data. However, since this method is effective at preserving previously learned information, its performance on time series data is uncertain. By focusing mainly on avoiding forgetting, it may fail to adequately capture new and evolving patterns in the data over time (Cossu et al., 2021; Pham et al., 2023). Additionally, these replay-based frameworks introduce significant complexity, as they must carefully determine which data samples to store in the buffer, when to update importance weights, and how to allocate memory and computational resources efficiently. These challenges highlight the need for a more streamlined and efficient approach to managing streaming time series data. Furthermore, the impact of catastrophic forgetting differs significantly in time series data and predictive modelling compared to non-temporal data and classification tasks, where class- and task-incremental learning are more common scenarios (Ao & Fayek, 2023; Pham et al., 2023). In such scenarios, the model’s objective is to learn new tasks or classes while ensuring that previously learned ones are not forgotten. On the other hand, when developing a lifelong learning model for streaming time series data, a key consideration emerges: to what extent should we prioritize preserving historical information, especially when that data may have evolved or changed over time (Cossu et al., 2021; Pham et al., 2023).
@@
Streaming time series forecasting faces significant challenges, including (1) Concept Drift Phenomenon, where underlying patterns or relationships within the data change over time, posing a substantial challenge for predictive modelling (Fekri et al., 2021). This issue extends beyond online learning models to impact batch learning approaches, as models initially proven accurate can quickly become obsolete as the data continues to evolve. (2) Temporal Irregularity, which is common in streaming data, especially in real-world applications where the data lacks a specific time frame. (3) Balanced Adaptation, where maintaining the right balance between historical and recent data is important to preventing biases towards outdated data while effectively adapting to the dynamic nature of streaming time series data. Therefore building a robust framework that can adeptly address these challenges is essential for maintaining model accuracy over time.
@@
This paper aims to overcome the restrictive boundaries of streaming time series processing by introducing a novel online learning model (ODEStream). Our focus lies on enabling the model to adapt to evolving data patterns without relying on a complex framework while maintaining good performance over time. Specifically, we investigate how the novel neural ordinary differential equation (ODE) methods can be leveraged for memory-free online forecasting for streaming time series data. The key contributions of this paper are as follows:
  • Unique Perspective: We redefine the context of continuous time series forecasting problem by introducing a novel perspective that prioritises learning the evolving dynamics and data distribution, rather than relying on the preservation of historical samples. This shift enables our model to adapt more effectively to changes over time, offering improved performance compared to traditional approaches.
@@
  • Utilising Neural ODEs: To the best of our knowledge, this work is the first to leverage neural Ordinary Differential Equation (ODE) models for continuous learning and online forecasting. Due to their ability to dynamically adapt and provide a continuous hidden state, ODEs offer a more flexible approach compared to traditional discrete-time models, especially in handling irregularly sampled data.
@@
  • Streamlined Buffer-Free Framework: We introduce ODEStream, a buffer-free continual learning framework designed to have a straightforward and efficient training process. Unlike existing frameworks that require complex decision-making for upcoming data; such as defining thresholds, trigger values, or managing buffered samples, ODEStream eliminates these requirements entirely. This streamlined design reduces overhead, simplifies training, and improves the model’s usability in real-world streaming scenarios.
@@
  • Comprehensive Evaluation: We evaluate ODEStream against state-of-the-art continual learning models and demonstrate that it effectively mitigates the impact of irregularities, adapts to data drift, and maintains stable performance over extended streaming periods.
@@ 2.2 Time Series Streaming Data
In general, online time series forecasting has primarily been studied using machine learning models (Jiménez-Herrera et al., 2023; Li et al., 2019; Chen et al., 2020; Pham et al., 2023; Melgar-García et al., 2023; Cossu et al., 2021; Fekri et al., 2021; Ao & Fayek, 2023) and statistical forecasting techniques (Alberg & Last, 2018), either using online-offline learning (Chen et al., 2020; Pham et al., 2023; Melgar-García et al., 2023), or by applying incremental learning (Alberg & Last, 2018; Li et al., 2019). However, using machine learning is not sufficient for multivariate real-world data, where the temporal dependence and pattern are complex and hard to discover. Recent works have focused on online time series forecasting using deep learning. The work in (Wambura et al., 2020) introduced a model named OFAT, which stands for One Sketch Fits All-Time Series. The method was designed for long-range forecasting in feature-evolving data streams, which involve data that changes over time and has varying features. The model employs deep neural networks to learn from a small sample of data streams and subsequently applies the learned patterns to new data streams. For that work, CNN is used to extract features from the input data, while the LSTM is used to predict future values of the output data.
In another work (Pham et al., 2023), the authors introduced a novel framework for online time series forecasting that combines the strengths of deep neural networks and Complementary Learning Systems (CLS). The framework, named FSNet, consists of two components: a slowly-learned backbone and a fast adapter. The backbone is responsible for learning long-term temporal dependencies, while the adapter is responsible for learning short-term changes. The two components are dynamically balanced to ensure that the model can adapt to both new and recurring patterns. When the data distribution changes, the adapter is updated to learn the new patterns. When the data distribution is stable, the backbone is updated to improve the accuracy of the prediction, while memory is used to store the historical data that the model has learned. Nevertheless, these works continue to face challenges when dealing with irregular samples.
More Recent studies have focused on the temporal distribution shift problem. In (Du et al., 2021), the concept of Temporal Covariate Shift (TCS) was introduced, and the AdaRNN framework was proposed. AdaRNN combines Temporal Distribution Characterisation (TDC) and Temporal Distribution Matching (TDM) to adaptively manage distribution changes. While the experimental results show that AdaRNN significantly improves forecasting accuracy, it relies heavily on correctly segmenting the data into diverse periods. Bai et al. (2022) introduced the DRAIN framework, which also focuses on adapting to changing data distributions over time by addressing the challenge of temporal domain generalization. DRAIN uses a Bayesian approach to jointly model data and neural network dynamics. It employs a recurrent graph generation scenario to encode and decode dynamic neural networks, capturing temporal drifts in both data and model parameters. A major drawback of the model is its requirement to encode and decode the entire network parameters, which adds significant complexity to the model, making it computationally intensive.
Another online learning algorithm for sequence data is Real Time Recurrent Learning (RTRL), which was recently revisited by (Irie et al., 2023) to improve its computational and memory requirements. The authors highlight RTRL’s conceptual advantages over Backpropagation Through Time (BPTT) along with the time and space complexity challenges. The paper explores approximation theories and tests RTRL in realistic settings and proposes a Real-Time Recurrent Actor-Critic (R2AC) framework that uses RTRL instead of standard Backpropagation Through Time (BPTT) for RNNs.
Recent work has also focused on online variational methods for state estimation in dynamic systems (Dowling et al., 2023; Campbell et al., 2021). In (Dowling et al., 2023), the authors introduce the exponential family variational Kalman filter (eVKF), an online recursive Bayesian method that infers latent neural trajectories while learning the underlying dynamical system. The eVKF effectively handles arbitrary likelihoods and offers a closed-form variational analogue to the Kalman filter’s prediction step, resulting in tighter ELBO bounds. Furthermore, in (Campbell et al., 2021) a method for online state estimation in state-space models (SSMs) was presented. The model operates by utilizing backward decompositions of the joint posterior distribution and Bellman-type recursions to maintain constant update costs.
@@ 2.3 Neural ODEs
Neural Ordinary Differential Equations (Neural ODEs) (Chen et al., 2018) represent a significant advancement in deep learning by integrating differential equations and neural networks. They provide a flexible framework for modelling complex dynamical systems, treating neural networks as continuous processes. Recent research in this field has made remarkable progress, showcasing the effectiveness of neural ODEs in various applications, especially for modelling irregularly sampled time series and partially observed data such as ODE-RNN, Latent-ODE, NJ-ODE and GRU-ODE models (Rubanova et al., 2019; Herrera et al., 2020; De Brouwer et al., 2019; Abushaqra et al., 2024). These models leverage the ODE’s capability to provide continuous hidden states, unlike the traditional discrete approach used in traditional sequence models such as RNN.
In recent years, several models based on Differential Equations (DE) have been introduced to address different issues, including modifying hidden trajectories (Kidger et al., 2020; Morrill et al., 2021; Schirmer et al., 2022) or reducing computational overhead (Habiba & Pearlmutter, 2020). However, these models have typically been evaluated and applied in batch learning and fixed datasets; hence, we do not consider them in this work. Despite the progress in applying neural ODEs to various domains, a critical gap exists in their application to streaming time series data. Existing works primarily focus on batch learning and fixed datasets, lacking exploration of the dynamic and evolving nature of streaming data. The intricacies of adapting neural ODEs to real-time processing of continuously changing data remain unexplored. Our research aims to bridge this gap by introducing ODEStream, a novel framework for continual learning in streaming time series, effectively utilising the strengths of neural ODEs in this challenging context. Moreover, while a recent work (Giannone et al., 2020) proposed real-time classification from event-camera streams using neural ODEs (INODE), its focus on high-frequency event data and real-time classification does not comprehensively address the broader spectrum of challenges associated with streaming time series prediction, which we aim to tackle in this study.
@@ 3 Preliminaries
In this section, we introduce the fundamental principles of ODEs-based models and ODE solvers, exploring the theory behind neural ODEs and their role in representing irregular time series.
@@ 4 Methodology
Problem statement: Let us consider a dynamic environment where streaming time series data is continuously generated as $X=\{x_{1},x_{2},\dots,x_{t},\dots\}$ with observations arriving in sequence over time. The goal is to develop a continual learning system for time series forecasting that adapts to the evolving nature of the data stream. Each observation $x_{t}$ represents the input at time $t$ for univariate data, while $x_{t}$ represents a vector of inputs at time $t$ for multivariate data, such as $x_{t}=[x_{t1},x_{t2},...,x_{td}]$, where $x_{td}$ is the value of the $d^{th}$ feature at time $t$. The objective is to predict the future values of the time series. The challenge lies in the continual learning setting, where the model must adapt and learn from the continuously incoming data without access to past observations.
Figure 1: ODEStream Framework: which leverages Neural ODEs to encode prior knowledge during the initial training phase. Subsequently, the learned model parameters are transferred and complemented by the temporal isolation layer for continuous online learning. In real-time, streaming data samples are used to predict future values, while the model simultaneously updates its parameters to adapt to newly observed data.
@@ 4.1 Neural ODEs for Continual Learning
Our goal is to design a training method for memory-free continual learning models. In the context of online learning, we aim for our model to keep learning and adapting as new data arrives without forgetting previously learned knowledge. Traditionally, complex models with buffers are used to store a subset of past data to tackle this issue. As discussed, time series continual learning studies have shown that memory-based approaches often achieve the best performance, particularly in classification tasks. These models continuously monitor incoming data and retrain on both old and new samples to maintain accuracy in dynamic environments, as demonstrated by Fekri et al. (2021). However, When these models are applied to continuous data streams, particularly those with gradually evolving distributions, they face significant challenges. The main issue is that the model only has one opportunity to train on each data sequence, unlike batch learning. Furthermore, in the presence of evolving temporal patterns, it remains unclear how the information stored in the buffer can contribute to future prediction tasks.
@@ 4.2 Overview of ODEStream
We proposed a streamlined model for continual learning and forecasting of time series, consisting of two main phases. $(i)$ an offline learning phase, commonly referred to as model warm-up, designed to initialise the model parameters. $(ii)$ online learning and forecasting, which generate predictions based on new sequence flow while adopting and learning from the new changes in the data. The first phase is instrumental in obtaining prior knowledge of the available time sequences. For this part, the model utilises variational autoencoders to capture the distribution, and temporal dependencies of historical data, and an ODE is employed to model evolving dynamics. The trained model is then transitioned to online learning, functioning as a dynamic adaptive network for continuous analysis of streaming sequences. In this primary phase, we leverage pre-trained Neural ODE weights. The model continuously adapts to changes in the data distribution by updating its understanding of the underlying dynamics in real time. As shown in Figure 1, It is noteworthy that the model framework lacks external memory for storing samples, and it does not include a decision node or threshold values to control the flow of the learning process. In the following sections, we provide a detailed description of the model framework.
4.3 Obtaining Prior Knowledge (Model warm-up)